OpenStudy (anonymous):

If u= 4 and v=5, and the vectors make angles of 60 degrees and 80 degrees with the x-axis respectively, find the component for of the sum of u and v.

OpenStudy (anonymous):

options are:. (2.87, 8.39) B. (1.13, - 1.46) C. (4.33, 6.92) D. (2.60, - 2.92) E. (3.19, 8.27)

OpenStudy (mathmate):

The given values of 4 and 5 are the magnitudes of the vectors u and v. t1=angle of u with x-axis = 60 deg. t2=angle of v with x-axis = 80 deg. The x-component of the sum = u cos(t1)+v cos(t2) = 4cos(60)+5cos(80)=2.86 The y-component of the sum = u sin(t1)+v sin(t2) = 4sin(60)+5sin(80) = 8.39 So look for your choice of answer,

OpenStudy (anonymous):

Thank you so much!

OpenStudy (mathmate):

You're welcome!